# Unfitted finite element modelling of surface-bulk viscous flows in animal cells

### v1

@@ 3.4. Solution strategy
Algorithm 1 exposes the need to carry information between consecutive discrete surfaces. In particular, the bulk unknowns $(\boldsymbol{u}_{h}^{n-1},p_{h}^{n-1})$, found in $\Omega_{h}^{n-1}$, have to be evaluated in $\Gamma_{h}^{n}=\partial\Omega^{n}$ (line 7); while the surface unknowns $\boldsymbol{U}_{h}^{n}$, $C_{h}^{n}$, found in $\Gamma_{h}^{n}$, have to be evaluated in $\Gamma_{h}^{n+1}$ (line 13). In order to ensure this, we follow a similar approach to the one detailed in, e.g., [lehrenfeld2018stabilized] for surface and [lehrenfeld2019eulerian, burman2022eulerian] for bulk problems. The main idea is to extend the discrete problems further away from the cells cut by the current surface/boundary, until they cover a region wide enough to contain the next surface (which is not known). This is achieved by formulating the unfitted surface FE problems in $\mathcal{T}_{h,n}^{\Gamma}=\left\{T\in\mathcal{T}_{h}:T\cap\{|\phi_{h}^{n}|<\delta^{n}\}\neq\emptyset\right\}$ and the unfitted bulk FE problem in $\mathcal{T}_{h,n}^{\Omega}=\left\{T\in\mathcal{T}_{h}:T\cap\{\phi_{h}^{n}<\delta^{n}\}\neq\emptyset\right\}$, where $\delta^{n}\propto\Delta{\rm t}\,\|\boldsymbol{U}_{h}^{n-1}\|_{\infty}$ is a width parameter. We refer to the above references for a complete description and discussion of this strategy.
@@ 4. Numerical experiments
Numerical implementation of the unfitted FE formulation in Section 3 was carried out using the Gridap.jl [Badia2020, Verdugo2022] FE software ecosystem, written in the Julia programming language. The code is available at the SurfaceBulkViscousFlows GitHub repository and supplemented with demonstrators of the numerical examples that follow. The unfitted FE tools are provided by GridapEmbedded.jl [gridapembedded_0_9_5]. It interfaces to the C++ algoim [algoim] library, via a Julia wrapper [algoimwrapper], to compute quadratures for domains implicitly defined by multivariate polynomials [saye2022high] and closest point projections on implicit surfaces [saye2014high]. We solve our linear systems with the parallel sparse direct solver MUMPS [MUMPS:1, MUMPS:2] available at the PETSc library v3.15.2 [petsc-web-page, petsc-user-ref, petsc-efficient]. Numerical results are postprocessed with ParaView v5.10.0 [AHRENS2005717].
@@ 4.1. Verification examples
We carry out two convergence tests to validate our code implementation. In the first experiment, we consider a steady-state 2D axisymmetric problem in a fixed geometry with known analytical solution. To this end, we transform the fluid sphere problem with two bulk phases of [happel2012low, Section 4-21] into a surface-bulk viscous flow problem. This leads to a coupled surface-bulk Stokes problem where (for this experiment only) the surface is material and, as a result, obeys the surface incompressibility condition. Using the method of manufactured solutions, we set the data of the problem such that the general solution for the surface and bulk phases, given by the stream functions in polar coordinates, reads
@@
The bulk phase is discretised with AgFEM, using the inf-sup stable pair $\mathcal{Q}_{2}\times\mathcal{P}_{1}$, as described in Section 3.2. In contrast, the surface phase is discretised with a consistent trace FE formulation for the surface Stokes problem [jankuhn2021error] using the inf-sup stable pair $\mathcal{Q}_{2}\times\mathcal{Q}_{1}$. Futher details of the implementation can be found in the code demonstrator of this example. As shown in Figure 5(d) optimal convergence rates of the iterative scheme are observed under uniform mesh refinements.
Figure 5. Verification examples. 2D axisymmetric surface-bulk fluid sphere problem adapted from [happel2012low, Section 4-21]: (a) Velocity magnitude and streamlines of the steady-state solution. (d) The rate of decay of the approximation error under uniform refinements obeys theoretical estimates [Badia2018Mixed, jankuhn2021error]. 2D axisymmetric dynamic surface: (b) Initial species concentration and velocity fields. Species concentration at the left sextant is 10% higher than elsewhere, setting the system out of equilibrium. (c) Equilibrium shape and solution in the comoving frame at $T=3.0$. (e) Decay of the experimental errors with uniform refinements. We observe rates of convergence close to quadratic order for all quantities. Note that the results are reflected across the axis of symmetry for illustration purposes.
@@
The initial density field of the molecular species (see Figure 5(b)) is set to $1.0$ everywhere, except on a sextant, where the value is increased to $1.1$. This initial bias implies the active tension on the surface is non-homogeneous, which cannot happen in a sphere. As a result, the system is forced into motion to find the equilibrium shape of Figure 5(c). In this process, the Marangoni effect increases the concentration on the sextant, while decreasing it on the other side. Meanwhile, the surface flows generate a reverse fountain flow in the bulk. At steady-state, the system has deformed into a bean-like shape, where the region of higher surface tension is flatter than the opposite side, the body travels at constant speed in the direction of the flows at the axis of symmetry, and the flow field in the comoving frame is tangent to the surface. This swimming mechanism mediated by Marangoni forces has been identified in the spontaneous self-propulsion of active colloids and droplets [zottl2016emergent].
@@ 4.2. Self-organised shape emergence
We now study the mechanochemical instabilities at the heart of the system formed by Equations (10)-(17), guiding spontaneous polarisation and pattern formation. Linear stability analysis [mietke2019minimal] of the rotationally symmetric model has shown the instabilities are regulated by the Péclet number Pe, relating active and diffusive surface transport, and the hydrodynamic length $L_{\eta}$, coupling the surface and bulk fluids.
@@
On the other hand, the hydrodynamic screening length $L_{\eta}$ categorises the stationary pattern. Symmetric or ring patterns appear when $L_{\eta}/R\ll 1$, while asymmetric or polar ones dominate when $L_{\eta}/R\gg 1$. In contrast to [mietke2019minimal, wittwer2023computational], we restrict ourselves to the latter case to reflect the common situation in animal cells; where typical 3D viscosities and geometrical scales of the cortex (thickness) and the cytoplasm (radius) yield rather large hydrodynamic lengths, with $L_{\eta}/R\approx 10^{3-4}$ [turlier2014furrow, valberg1987magnetic]. This type of polar instabilities, at large hydrodynamic lengths, have been suggested as a mechanism for sustained unidirectional cell motility in three-dimensional environments [hawkins2011spontaneous].
@@
Using real spherical harmonics $Y_{lm}(\theta,\varphi)$, linear stability analysis of a homogeneous stationary state leads to a critical Péclet number [mietke2019minimal]
@@ 4.3. Relaxation dynamics
We continue analysing the mechanochemical instabilities of Equations (10)-(17) focusing on a different driving force: curvature inhomogeneities [mietke2019self]. The linear stability analysis of Section 4.2 identifies the sphere as a stable shape for $\mbox{{Pe}}\!<\!\mbox{{Pe}}_{\rm cr}^{l}$, $l\geq 1$, in which diffusion, dominating over advection, homogenenises the concentration of the molecular species. On the other hand, recalling Equation (9), the decomposition of the active forces exposes the capability of mean curvature gradients alone to generate the Marangoni effect. Obviously, these cannot be balanced out by diffusive transport. Hence, in the diffusion-dominated regime, the system is brought to relax its shape towards a sphere.
@@
Bounding box $\mathcal{B}$ $[-2.01,2.01]^{3}$ $[-2.25,1.25]\times[-1.75,1.75]^{2}$ $[-1.1,1.1]^{3}$
@@
Figure 10. Relaxation dynamics. Marangoni effects generated by mean curvature gradients are balanced by evolving the shape towards a sphere. Relaxation of a popcorn, a pear and a torus. (a)-(d) and (i)-(l) Surface concentration of stress-regulating molecular species. (e)-(h) and (m)-(p) Bulk flow field.
@@ 4.4. Bi- and uni-lateral cytokinesis
We now showcase the capabilities of the unfitted FE framework to deal with constricted deformations. The goal is to simulate cytokinesis, i.e., the physical process at the end of cell division that partitions the cell into two daughter cells [green2012cytokinesis]. In animal cells, cytokinesis is achieved by an actomyosin furrow forming at the equator that constricts the cell until scission. Meanwhile, cortical tension at the poles resists the increasing cell pressure [sedzinski2011polar]. As reviewed in [da2022viscous], both early hypotheses and recent experimental and modelling studies support the idea that cell division is driven by a gradient of surface tension directed toward the division axis. Moreover, the dominant mechanical contributor is cortical tension, rather than active or viscous torque contributions [turlier2014furrow, da2022viscous]. The mathematical model of Equations (10)-(17) is thus adequate for this type of cellular process.
We first focus on the stereotypical mode of cell division, i.e., symmetric cytokinesis, which cuts the cell into two equal-sized daughter cells. This type of division is achieved by a circumferential cleavage furrow that forms and develops a ring constriction with two cortical poles under tensile stress at both sides. We will then refer to this configuration as bilateral cytokinesis. The starting point is a spherical cell of radius $R_{0}$. Experimental measurements have shown that the mitotic apparatus of animal cells generates and sustains a Gaussian-like band of myosin overactivity to position the furrow and guide the constriction at the cell equator [rappaport1996cytokinesis, bement2005microtubule]. Following [turlier2014furrow], we model the overactivity at the contractile ring by spatially modulating the active surface tension $\mathbf{N}_{\Gamma}^{\rm act}$ in Equation (8) through the myosin activity field $\xi$, defined as
@@
where $\xi_{0}$ represents a basal level of activity, controlling active tension at the poles, $\delta\xi$ is the overactivity, expanding along the axis of rotational symmetry $x_{\alpha}$ over a typical width $\omega$. Figure 11 and 12 show the evolution of myosin concentration at the surface and the flows in the cytoplasm for the 2D axisymmetric and 3D models. We prescribe $R_{0}\!=\!0.46$, $\omega\!=\!0.1$, $\xi_{0}\!=\!1$ and $\delta\xi\!=\!10$; the rest of simulation parameters are listed in Table 2. We observe the furrow constriction appearing due to overactivity and being reinforced by flows of myosin from the poles to the equator. Flow patterns in the cytoplasm, with two toroidal vortices forming at each side of the equator, agree with experimental observations [hiramoto1958quantitative]. Figure 13 represents the evolution of the radius of the contractile ring $r_{f}$ and the distance between the poles $d_{p}$, relative to the initial cell radius $R_{0}$. We observe tangent flows at the poles when $d_{p}$ plateaus.
@@
Bounding box $\mathcal{B}$ $[-0.75,0.75]\times[0.0,0.75]$ $[-0.75,0.75]^{3}$
@@
Our last experiment targets the non-rotationally symmetric mode of cell division known as unilateral cytokinesis [sugioka2022symmetry]. In the zygote of certain animal species, including jellyfish, corals and comb jellies (cnidaria and ctenophora, in general) [rappaport1963experimental], but also in some epithelial tissues [herszterg2013interplay], the cleavage furrow ingresses from only one side of the cell. This implies the formation of a single extensile pole at the opposite side of the furrow. Figure 14 describes the dynamics of myosin concentration on the surface and bulk viscous flows during a unilateral cytokinesis that takes the same parameters as the bilateral case, except that $\xi$ in Equation (28) only acts in half the equator. In contrast to the bilateral case, only one toroidal vortex forms around the ingression furrow.
@@
We finally observe that both 2D axisymmetric and 3D models fail to complete cell pinching: In the 2D axisymmetric case, the axisymmetry boundary condition at the bulk phase (zero vertical displacement) is not compatible with the downwards displacement of the cleavage furrow. In the 3D cases, we observe shape instabilities arising when approximating pinch off. Both issues are to be investigated in future work.
@@ 5. Conclusions
Morphogenetic processes are frequently characterised by very intricate three-dimensional and shape-evolving fluid dynamics, such as the cortex-cytoplasm interactions that concern this work. FE analysis is one of the few computational approaches suitable for modelling this level of complexity, making it relatively easy to solve multiphysics PDE problems, with couplings between biochemical and mechanical signaling and between different topological dimensions. However, problems with moving boundaries and interfaces have been a major challenge in the FE community. This is due –in part– to the difficulty with FE methods to strike a good balance between coping with large deformations and accurately resolving free surfaces and internal interfaces [Donea2004].
@@
Nowadays, a new generation of unfitted FE methods is challenging the state-of-the-art in FEs for moving boundaries and interfaces [de2023stability, Olshanskii2017, Hansbo2020]. They rely on a Eulerian description of motion and a sharp-interface representation of the geometry. By providing a robust and accurate way to track moving surfaces in a fixed computational grid, they are questioning the classical trade-off between Lagrangian and Eulerian FE methods [Donea2004]. In this work, we specialise this emerging class of unfitted FE technologies to model surface-bulk viscous flows in animal cells. To this end, we have formulated a novel partitioned FE method that combines the aggregated FEM [Badia2018Mixed] for the bulk PDEs with trace FEM [jankuhn2021trace] for the surface PDEs. We have implemented the numerical model in Gridap.jl [Badia2020, Verdugo2022], an advanced Julia FE software ecosystem, leveraging high-order algorithms to compute quadrature rules and closest-point projections from the algoim library [saye2014high, saye2022high]. Our numerical experiments consider a minimal model of cellular symmetry breaking [mietke2019minimal]. They illustrate the capacity of the numerical framework to simulate nontrivial 3D dynamics, with very large distortions and topological changes, and the potential to address applications in animal morphogenesis that have been barely object of numerical modelling.
@@
Our unfitted methodology offers several promising avenues for future development to enhance both its fidelity and applicability to address increasingly complex morphogenetic phenomena. A natural extension would involve a more faithful representation of the actomyosin cortex as a thin shell, incorporating membrane (e.g., tension) and torque stress resultants previously derived in [da2022viscous]. While membrane tension resultants can be readily integrated into the current model, capturing surface torques—whether of active [salbreux2017mechanics], viscous [da2022viscous], or elastic origin (such as plasma membrane bending [torres2019modelling])—requires second-order derivatives of the unknown fields. This necessity motivates an extension of our framework toward $\mathcal{H}^{2}$-conforming approximations [gfrerer2021ac].
@@
Beyond morphogenetic deformations, our approach could be extended to study 3D cell motility in viscous environments [poincloux2011contractility, hawkins2011spontaneous, farutin2019crawling], including key additional mechanisms such as the membrane-to-cortex mechanical coupling [de2023cell], and the mechanochemical interaction with regulatory proteins [gross2017active]. In particular, mutually inhibitory signaling circuits—often modeled by nonlinear reaction-diffusion systems [goehring2011polarization, gross2019guiding, de2024long]—play a key role in symmetry breaking and polarity establishment.
@@
Mechanochemical coupling between cortical mechanics, surface-localized active proteins, and bulk-distributed passive species [howard2011turing] generally involves surface-bulk exchange of molecular components. Such coupling has been shown to underlie intracellular pattern formation [halatek2018self, brauns2021bulk, francis2024spatial]. Our unfitted framework can be easily adapted to simulate such processes, by solving advection-reaction-diffusion equations both on the cortex and in the cytoplasmic bulk, while preserving a sharp interface description. This feature provides a significant advantage over diffuse-interface or phase-field approaches [marth2014signaling, camley2017crawling, aland2023phase], enabling a more accurate resolution of the interface, and avoids alternative approximations using projection methods [burkart2024dimensionality].
@@
Further developments could incorporate the polar and nematic ordering of cortical actin filaments [salbreux2009hydrodynamics, salbreux2022theory], which are known to interact with gradients in cortical flow during cell division [reymann2016cortical, Mirza_2024]. Our unfitted framework could again prove advantageous in this context, simplifying the otherwise technical body-fitted FE formulation and implementation of tensorial fields on moving curved surfaces [nestler2019finite, torres2020approximation, nitschke2025active].
@@
As for modelling the bulk cytoplasm, we have treated the medium as a passive viscous fluid. However, the cytoplasm is a heterogeneous, crowded medium composed of a cytoskeletal meshwork permeated by cytosol and embedding various membrane-bound organelles such as the nucleus, endoplasmic reticulum, Golgi apparatus, and mitochondria. Capturing such complexity would require, at least, a two-phase poroviscous or poroelastic model [moeendarbary2013cytoplasm, mogilner2018intracellular], potentially coupled to elastic-like inclusions [royer2019quasi, Liao2024].
@@
Finally, a natural direction for generalization lies in the extension to multicellular systems, following surface-based approaches as in [maitre2016asymmetric, torres2022interacting, firmin2024mechanics]. This would pave the way for studying surface-bulk interactions in early animal embryogenesis, a domain in which computational modelling remains scarce—particularly using unfitted finite element methods. Overall, the intersection between finite element technologies and animal morphogenesis offers a wide frontier of methodological and biological challenges, and we anticipate this cross-disciplinary interaction to remain fertile and active for years to come.

### v2

@@ 3.4. Solution strategy
Algorithm 1 exposes the need to carry information between consecutive discrete surfaces. In particular, the bulk unknowns $(\boldsymbol{u}_{h}^{n-1},p_{h}^{n-1})$, found in $\Omega_{h}^{n-1}$, have to be evaluated in $\Gamma_{h}^{n}=\partial\Omega^{n}$ (line 7); while the surface unknowns $\boldsymbol{U}_{h}^{n}$, $C_{h}^{n}$, found in $\Gamma_{h}^{n}$, have to be evaluated in $\Gamma_{h}^{n+1}$ (line 13). In order to ensure this, we follow a similar approach to the one detailed in, e.g., [89] for surface and [90, 31] for bulk problems. The main idea is to extend the discrete problems further away from the cells cut by the current surface/boundary, until they cover a region wide enough to contain the next surface (which is not known). This is achieved by formulating the unfitted surface FE problems in $\mathcal{T}_{h,n}^{\Gamma}=\left\{T\in\mathcal{T}_{h}:T\cap\{|\phi_{h}^{n}|<\delta^{n}\}\neq\emptyset\right\}$ and the unfitted bulk FE problem in $\mathcal{T}_{h,n}^{\Omega}=\left\{T\in\mathcal{T}_{h}:T\cap\{\phi_{h}^{n}<\delta^{n}\}\neq\emptyset\right\}$, where $\delta^{n}\propto\Delta{\rm t}\,\|\boldsymbol{U}_{h}^{n-1}\|_{\infty}$ is a width parameter. We refer to the above references for a complete description and discussion of this strategy.
Regarding the stability of the coupling scheme, numerical experiments in Section 4 adopt rather small time steps $\Delta{\rm t}\approx 10^{-4}-10^{-5}$ to correctly capture fast shape changes and, in particular, sharp nonlinear transitions of self-organised shape emergence, as in Figure 8. Hence, we reasonably assume stability of the scheme for such small time increments. For larger time steps, a rigorous stability analysis must be carried out.
@@ 4. Numerical experiments
Numerical implementation of the unfitted FE formulation in Section 3 was carried out using the Gridap.jl [10, 150] FE software ecosystem, written in the Julia programming language. The code is available at the SurfaceBulkViscousFlows GitHub repository and supplemented with demonstrators of the numerical examples that follow. The unfitted FE tools are provided by GridapEmbedded.jl [152]. It interfaces to the C++ algoim [135] library, via a Julia wrapper [107], to compute quadratures for domains implicitly defined by multivariate polynomials [134] and closest point projections on implicit surfaces [133]. We solve our linear systems with the parallel sparse direct solver MUMPS [4, 3] available at the PETSc library v3.15.2 [12, 11, 13]. Numerical results are postprocessed with ParaView v5.10.0 [1].
@@ 4.1. Verification examples
We carry out two convergence tests to validate our code implementation. In the first experiment, we consider a steady-state 2D axisymmetric problem in a fixed geometry with known analytical solution. To this end, we transform the fluid sphere problem with two bulk phases of [70, Section 4-21] into a surface-bulk viscous flow problem. This leads to a coupled surface-bulk Stokes problem where (for this experiment only) the surface is material and, as a result, obeys the surface incompressibility condition. Using the method of manufactured solutions, we set the data of the problem such that the general solution for the surface and bulk phases, given by the stream functions in polar coordinates, reads
@@
The numerical approximation adopts FE pairs with proven inf-sup stability: The bulk phase is discretised with AgFEM, using the inf-sup stable pair $\mathcal{Q}_{2}\times\mathcal{P}_{1}$ [7], as described in Section 3.2. In contrast, the surface phase is discretised with a consistent trace FE formulation for the surface Stokes problem [80] using the inf-sup stable pair $\mathcal{Q}_{2}\times\mathcal{Q}_{1}$. Futher details of the implementation can be found in the code demonstrator of this example. As shown in Figure 6(d) optimal convergence rates of the iterative scheme are observed under uniform mesh refinements.
Figure 6. Verification examples. 2D axisymmetric surface-bulk fluid sphere problem adapted from [70, Section 4-21]: (a) Velocity magnitude and streamlines of the steady-state solution. (d) The rate of decay of the approximation error under uniform refinements obeys theoretical estimates [7, 80]. 2D axisymmetric dynamic surface: (b) Initial species concentration and velocity fields. Species concentration at the left sextant is 10% higher than elsewhere, setting the system out of equilibrium. (c) Equilibrium shape and solution in the comoving frame at $T=3.0$. (e) Decay of the experimental errors with uniform refinements. We observe rates of convergence close to quadratic order for all quantities. Note that the results are reflected across the axis of symmetry for illustration purposes.
@@
The initial density field of the molecular species (see Figure 6(b)) is set to $1.0$ everywhere, except on a sextant, where the value is increased to $1.1$. This initial bias implies the active tension on the surface is non-homogeneous, which cannot happen in a sphere. As a result, the system is forced into motion to find the equilibrium shape of Figure 6(c). In this process, the Marangoni effect increases the concentration on the sextant, while decreasing it on the other side. Meanwhile, the surface flows generate a reverse fountain flow in the bulk. At steady-state, the system has deformed into a bean-like shape, where the region of higher surface tension is flatter than the opposite side, the body travels at constant speed in the direction of the flows at the axis of symmetry, and the flow field in the comoving frame is tangent to the surface. This swimming mechanism mediated by Marangoni forces has been identified in the spontaneous self-propulsion of active colloids and droplets [161].
@@ 4.2. Self-organised shape emergence
We now study the mechanochemical instabilities at the heart of the system formed by Equations (10)-(17), guiding spontaneous polarisation and pattern formation. Linear stability analysis [98] of the rotationally symmetric model has shown the instabilities are regulated by the Péclet number Pe, relating active and diffusive surface transport, and the hydrodynamic length $L_{\eta}$, coupling the surface and bulk fluids.
@@
On the other hand, the hydrodynamic screening length $L_{\eta}$ categorises the stationary pattern. Symmetric or ring patterns appear when $L_{\eta}/R\ll 1$, while asymmetric or polar ones dominate when $L_{\eta}/R\gg 1$. In contrast to [98, 157], we restrict ourselves to the latter case to reflect the common situation in animal cells; where typical 3D viscosities and geometrical scales of the cortex (thickness) and the cytoplasm (radius) yield rather large hydrodynamic lengths, with $L_{\eta}/R\approx 10^{3-5}$, see Table 1. This type of polar instabilities, at large hydrodynamic lengths, have been suggested as a mechanism for sustained unidirectional cell motility in three-dimensional environments [71].
Notation Quantity Experimental values References
[…]
$R$ cell radius (large cells) 5010050-100 µm [76, 140, 92]
$e$ cortical thickness 0.220.2-2 µm [37]
Table 1. Self-organised shape emergence. Experimental values reported in the literature for large animal cells leading to the estimate $L_{\eta}/R=\lambda_{\Gamma}/\mu_{\Omega}R=\lambda_{\Gamma}^{\rm 3D}e/\mu_{\Omega}R\approx 10^{3-5}$.
@@
Using real spherical harmonics $Y_{lm}(\theta,\varphi)$, linear stability analysis of a homogeneous stationary state leads to a critical Péclet number [98]
@@ 4.3. Relaxation dynamics
We continue analysing the mechanochemical instabilities of Equations (10)-(17) focusing on a different driving force: curvature inhomogeneities [99]. The linear stability analysis of Section 4.2 identifies the sphere as a stable shape for $\mbox{{Pe}}\!<\!\mbox{{Pe}}_{\rm cr}^{l}$, $l\geq 1$, in which diffusion, dominating over advection, homogenenises the concentration of the molecular species. On the other hand, recalling Equation (9), the decomposition of the active forces exposes the capability of mean curvature gradients alone to generate the Marangoni effect. Obviously, these cannot be balanced out by diffusive transport. Hence, in the diffusion-dominated regime, the system is brought to relax its shape towards a sphere.
@@
Bounding box $\mathcal{B}$ $01,2.01]^{3} [2.2$$25,1.25]\times[-1.75,1.75]^{2}$ $1.1,1.1]^{3}$
@@
Figure 11. Relaxation dynamics. Marangoni effects generated by mean curvature gradients are balanced by evolving the shape towards a sphere. Relaxation of a popcorn, a pear and a torus. (a)-(d), (i)-(l) and (q)-(t) Surface concentration of stress-regulating molecular species. (e)-(h), (m)-(p) and (u)-(x) Bulk flow field.
@@ 4.4. Bi- and uni-lateral cytokinesis
We now showcase the capabilities of the unfitted FE framework to deal with constricted deformations. The goal is to simulate cytokinesis, i.e., the physical process at the end of cell division that partitions the cell into two daughter cells [63]. In animal cells, cytokinesis is achieved by an actomyosin furrow forming at the equator that constricts the cell until scission. Meanwhile, cortical tension at the poles resists the increasing cell pressure [139]. As reviewed in [40], both early hypotheses and recent experimental and modelling studies support the idea that cell division is driven by a gradient of surface tension directed toward the division axis. Moreover, the dominant mechanical contributor is cortical tension, rather than active or viscous torque contributions [148, 40]. The mathematical model of Equations (10)-(17) is thus adequate for this type of cellular process.
We first focus on the stereotypical mode of cell division, i.e., symmetric cytokinesis, which cuts the cell into two equal-sized daughter cells. This type of division is achieved by a circumferential cleavage furrow that forms and develops a ring constriction with two cortical poles under tensile stress at both sides. We will then refer to this configuration as bilateral cytokinesis. The starting point is a spherical cell of radius $R_{0}$. Experimental measurements have shown that the mitotic apparatus of animal cells generates and sustains a Gaussian-like band of myosin overactivity to position the furrow and guide the constriction at the cell equator [124, 17]. Following [148], we model the overactivity at the contractile ring by spatially modulating the active surface tension $\mathbf{N}_{\Gamma}^{\rm act}$ in Equation (8) through the myosin activity field $\xi$, defined as
@@
where $\xi_{0}$ represents a basal level of activity, controlling active tension at the poles, $\delta\xi$ is the overactivity, expanding along the axis of rotational symmetry $x_{\alpha}$ over a typical width $\omega$. Figure 12 and 13 show the evolution of myosin concentration at the surface and the flows in the cytoplasm for the 2D axisymmetric and 3D models. We prescribe $R_{0}\!=\!0.46$, $\omega\!=\!0.1$, $\xi_{0}\!=\!1$ and $\delta\xi\!=\!10$; the rest of simulation parameters are listed in Table 3. We observe the furrow constriction appearing due to overactivity and being reinforced by flows of myosin from the poles to the equator. Flow patterns in the cytoplasm, with two toroidal vortices forming at each side of the equator, agree with experimental observations [76]. Figure 14 represents the evolution of the radius of the contractile ring $r_{f}$ and the distance between the poles $d_{p}$, relative to the initial cell radius $R_{0}$. We observe tangent flows at the poles when $d_{p}$ plateaus.
@@
Bounding box $\mathcal{B}$ $0.75]\times[0.0,0.75] [0.75,0.75$$0.75]^{3}$
@@
Our last experiment targets the non-rotationally symmetric mode of cell division known as unilateral cytokinesis [142]. In the zygote of certain animal species, including jellyfish, corals and comb jellies (cnidaria and ctenophora, in general) [123], but also in some epithelial tissues [75], the cleavage furrow ingresses from only one side of the cell. This implies the formation of a single extensile pole at the opposite side of the furrow. Figure 15 describes the dynamics of myosin concentration on the surface and bulk viscous flows during a unilateral cytokinesis that takes the same parameters as the bilateral case, except that $\xi$ in Equation (28) only acts in half the equator. In contrast to the bilateral case, only one toroidal vortex forms around the ingression furrow.
@@
We finally observe that both 2D axisymmetric and 3D models fail to complete cell pinching: In the 2D axisymmetric case, the axisymmetry boundary condition at the bulk phase (zero vertical displacement) is not compatible with the downwards displacement of the cleavage furrow. In the 3D cases, we observe shape instabilities arising when approximating pinch off. On the other hand, we identified that alternative models for cytokinesis with stress-dependent myosin detachment rates or stress-dependent surface shear viscosities [24] require to solve the transport problem with a mass-conserving scheme, which we have not included in the scope of this work. All these issues will be investigated in future work.
@@ 5. Conclusions
Morphogenetic processes are frequently characterised by very intricate three-dimensional and shape-evolving fluid dynamics, such as the cortex-cytoplasm interactions that concern this work. FE analysis is one of the few computational approaches suitable for modelling this level of complexity, making it relatively easy to solve multiphysics PDE problems, with couplings between biochemical and mechanical signaling and between different topological dimensions. However, problems with moving boundaries and interfaces have been a major challenge in the FE community. This is due –in part– to the difficulty with FE methods to strike a good balance between coping with large deformations and accurately resolving free surfaces and internal interfaces [49].
@@
Nowadays, a new generation of unfitted FE methods is challenging the state-of-the-art in FEs for moving boundaries and interfaces [44, 113, 68]. They rely on a Eulerian description of motion and a sharp-interface representation of the geometry. By providing a robust and accurate way to track moving surfaces in a fixed computational grid, they are questioning the classical trade-off between Lagrangian and Eulerian FE methods [49]. In this work, we specialise this emerging class of unfitted FE technologies to model surface-bulk viscous flows in animal cells. To this end, we have formulated a novel partitioned FE method that combines the aggregated FEM [7] for the bulk PDEs with trace FEM [81] for the surface PDEs. We have implemented the numerical model in Gridap.jl [10, 150], an advanced Julia FE software ecosystem, leveraging high-order algorithms to compute quadrature rules and closest-point projections from the algoim library [133, 134]. Our numerical experiments consider a minimal model of cellular symmetry breaking [98]. They illustrate the capacity of the numerical framework to simulate nontrivial 3D dynamics, with very large distortions and topological changes, and the potential to address applications in animal morphogenesis that have been barely object of numerical modelling.
@@
Our unfitted methodology offers several promising avenues for future development to enhance both its fidelity and applicability to address increasingly complex morphogenetic phenomena. A natural extension would involve a more faithful representation of the actomyosin cortex as a thin shell, incorporating membrane (e.g., tension) and torque stress resultants previously derived in [40]. While membrane tension resultants can be readily integrated into the current model, capturing surface torques—whether of active [130], viscous [40], or elastic origin (such as plasma membrane bending [146])—requires second-order derivatives of the unknown fields. This necessity motivates an extension of our framework toward $\mathcal{H}^{2}$-conforming approximations [60].
@@
Beyond morphogenetic deformations, our approach could be extended to study 3D cell motility in viscous environments [118, 71, 53], including key additional mechanisms such as the membrane-to-cortex mechanical coupling [42], and the mechanochemical interaction with regulatory proteins [65]. In particular, mutually inhibitory signaling circuits—often modeled by nonlinear reaction-diffusion systems [61, 64, 41]—play a key role in symmetry breaking and polarity establishment.
@@
Mechanochemical coupling between cortical mechanics, surface-localized active proteins, and bulk-distributed passive species [78] generally involves surface-bulk exchange of molecular components. Such coupling has been shown to underlie intracellular pattern formation [67, 27, 57]. Our unfitted framework could address these scenarios with mass-conserving or, more generally, structure-preserving schemes [39] for the robust solution of advection-reaction-diffusion equations both on the cortex and in the cytoplasmic bulk, while preserving a sharp interface description. This feature provides a significant advantage over diffuse-interface or phase-field approaches [96, 34, 2], enabling a more accurate resolution of the interface, and avoids alternative approximations using projection methods [29].
@@
Further developments could incorporate the polar and nematic ordering of cortical actin filaments [131, 129], which are known to interact with gradients in cortical flow during cell division [125, 144]. Our unfitted framework could again prove advantageous in this context, simplifying the otherwise technical body-fitted FE formulation and implementation of tensorial fields on moving curved surfaces [108, 147, 110].
@@
As for modelling the bulk cytoplasm, we have treated the medium as a passive viscous fluid. However, the cytoplasm is a heterogeneous, crowded medium composed of a cytoskeletal meshwork permeated by cytosol and embedding various membrane-bound organelles such as the nucleus, endoplasmic reticulum, Golgi apparatus, and mitochondria. Capturing such complexity would require, at least, a two-phase poroviscous or poroelastic model [102, 103], potentially coupled to elastic-like inclusions [126, 91]. Stable and geometrically robust unfitted FE formulations for these type of problems are still rather limited and object of ongoing research [160, 22].
@@
Finally, a natural direction for generalization lies in the extension to multicellular systems, following surface-based approaches as in [94, 145, 55]. This would pave the way for studying surface-bulk interactions in early animal embryogenesis, a domain in which computational modelling remains scarce—particularly using unfitted finite element methods. Extending our FE framework for evolving multicellular systems faces several challenges that include, but are not limited to, the capability of handling full cell division and pinch-off and an automatic approach to label cells and their free and internal surfaces. Overall, the intersection between finite element technologies and animal morphogenesis offers a wide frontier of methodological and biological challenges, and we anticipate this cross-disciplinary interaction to remain fertile and active for years to come.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper
CRediT authorship contribution statement
Eric Neiva: Conceptualization, Funding acquisition, Formal analysis, Investigation, Methodology, Software, Visualization, Writing - original draft preparation, review and editing. Hervé Turlier: Conceptualization, Funding acquisition, Methodology, Resources, Supervision, Validation, Writing - review and editing.